graph the linear equation
x+4=0

Answer by jim_thompson5910(35256)

x+4=0

x+4-4=0-4

x = -4

So this is the vertical line at the tick -4 on the x axis


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim